# LLM Collaboration with Multi-Agent Reinforcement Learning

### v2

@@ 5.2 Baselines
We adopt the fixed and fine-tuned single model, along with three multi-agent methods built on fixed base models, as our baselines. For the single-agent setting, the Qwen2.5-Coder-3B model generates a Python function based on the problem description in the prompt, with the function name specified in entry_point. We also fine-tune this model on the training set to adapt it to the task. In the multi-agent setting, two Qwen2.5-Coder-3B models serve as agents: one generates an auxiliary function, and the other produces the main function. To minimize the influence of prompts on our comparison, we keep the problem description fixed and only add minimal coordination instructions. Specifically, in the naive concatenation scheme, agents are simply informed of their roles and generate outputs in parallel without communication. The sequential pipeline introduces one-way communication, allowing the main agent to respond based on the auxiliary agent’s output. The one-round discussion baseline enables bidirectional communication: agents first receive the same prompts as in naive concatenation, then the prompts are augmented with the other’s first-turn response in the second turn.
5.3 Results
We optimize the interaction between 2 agents with MAGRPO in both single-turn and multi-turn settings, i.e., MAGRPO-ST and MAGRPO-MT. To reduce prompt-induced variance, agents are informed only of the problem description and their roles, using the same initial prompt as in naive concatenation. In MAGRPO-MT, the agents’ previous responses are given to an external Claude-Sonnet-4 model, which provides each agent with feedback comprising functionality analysis, error detection, and revision suggestions.
The performance of MAGRPO-ST and MAGRPO-MT on HE is shown in Figures 2(a) and 2(b). Although MAGRPO-ST improves the syntactical correctness and develops valid cooperation, its test pass rate does not show much progress. As for MAGRPO-MT, agents are initially overwhelmed by the external model’s feedback, resulting in even lower initial returns. They gradually adopt the suggestions and improve their returns. However, the improvement in test pass rate is still limited due to noisy entries in the dataset and unreliable feedback. This reflects the complexity and delicacy of coder coordination, where the main agent must accurately infer the functionality of auxiliary modules and trust their correctness without direct communication. As shown in Figures 2(c) and 2(d), MAGRPO-ST and MAGRPO-MT achieve higher overall returns and lower variances when trained on CHE. In MAGRPO-MT, although agents initially struggle to interpret the feedback, the normalized returns gradually increase and eventually surpass those of single-turn training. This indicates that, when trained on a dataset with well-defined cooperative structures and guided by reliable suggestions, agents can learn to incorporate feedback effectively and improve the quality of their responses.
Method Speed Pass@1 Pass@3 Pass@5 Pass@10
HE CHE HE CHE HE CHE HE CHE HE CHE
Fixed Single Model 113.8 125.6 38.7 50.0 54.8 56.3 61.2 62.5 67.7 75.0
Fine-Tuned Single Model 114.9 124.8 45.1 62.5 64.5 68.8 67.7 75.0 70.9 81.0
Naive Concatenation 194.9 189.4 42.5 40.1 45.2 43.8 50.6 56.3 64.5 59.2
Sequential Pipeline 99.6 97.4 53.4 57.2 54.8 73.6 62.3 82.7 71.2 86.2
One-Round Discussion 82.5 78.3 41.2 40.9 51.6 52.9 61.1 60.3 70.8 75.3
MAGRPO-ST (Ours) 190.0 192.4 54.8 68.2 55.6 71.4 58.1 75.4 71.6 81.2
MAGRPO-MT (Ours) 95.2 97.3 67.9 73.2 71.0 75.0 80.6 81.3 90.3 87.5
@@ 5.3 Results
Table 1: Performance of MAGRPO and baselines on speed (tokens/s) and pass@k. Results are averaged over 10 runs; bold indicates the best performance for each metric on each dataset.
@@
Table 1 presents a performance comparison between MAGRPO and baselines on HE and CHE. Speed is measured in tokens per second on a GeForce RTX 5090, and pass@k is shown in percentage. By GRPO fine-tuning, the performance of Qwen2.5-Coder-3B model only improves slightly as the logic of test problems differs substantially from that in the training set. Although the naive concatenation method exhibits high generation speed, it has lower test pass rates than a single model, as the main agent may rely on incorrect assumptions about the auxiliary function. In the sequential pipeline, the main agent can access the auxiliary function and compensate for its weaknesses, improving robustness but at the cost of slower inference. The one-round discussion method involves more communication between agents, but its effectiveness remains limited due to the potential misaligned cross-adaptation issue. MAGRPO-ST incorporates diverse cooperation schemes that improve robustness, thereby obtaining higher unit pass rates. In MAGRPO-MT, an external model can provide additional feedback that further strengthens cooperation. As a result, MAGRPO-MT achieves the highest pass rates and total return across most pass@k metrics.
5.4 Cooperation Schemes and Learning Modes
@@
MAGRPO identifies various cooperation schemes. For example, the auxiliary function handles the core logic, while the main agent adds backup logic or decorations to improve the overall solution. Alternatively, the main agent may act as a coordinator, decomposing the problem and assigning subtasks to the auxiliary agent. In addition, the auxiliary function may also serve as a strategy filter, guiding the main agent to generate code for specific cases. While coordinator and strategy-filter schemes can improve inference efficiency, they are more prone to syntax and logical errors. With limited cooperation-oriented training data, the main agent typically resorts to more conservative roles, i.e., fallback or decoration. Note that these cooperation schemes emerge during training under a relatively simple joint reward, more refined design patterns could be found when training agents to develop large-scale software.
(a) Self-Evolving
(b) Guided-Learning
Figure 3: MAGRPO-MT rewards in 2 learning modes: (a) Turn 1 (blue); (b) Turn 2 (red).
Agents can learn to cooperate through various modes in the multi-turn setting. Figure 3(a) demonstrates a self-evolving mode, where agents primarily evolve with the tasks themselves without external feedback. At the beginning, the second-turn rewards (red) are lower than the first-turn rewards (blue), since agents struggle to incorporate previous responses effectively. Both curves gradually improve as agents develop cooperative behaviors, but the performance of the second turn is consistently similar to the first turn, suggesting that only providing the previous responses is ineffective or hardly interpreted by the agents. Figure 3(b) illustrates guided-learning mode, where LLMs leverage external feedback to improve performance. When using Claude-Sonnet-4 to provide concrete suggestions (e.g., code edits), the performance of the second turn (red) exceeds first turn (blue), and both outperform those in the self-evolving, indicating that appropriate guidance helps agents to refine the response. Due to the computational constraints, most models used in our setup have around 3B parameters and struggle to interpret vague feedback. We hypothesize that larger models with better reasoning abilities could benefit from more implicit guidance.
6 Conclusion
@@
In this paper, we model LLM collaboration as a cooperative MARL problem and formalize it as a Dec-POMDP. We introduce the MAGRPO algorithm to optimize their cooperation with aligned rewards. Our experiments on coding collaboration show that MAGRPO enables agents to generate higher-quality solutions more efficiently through effective coordination. This work highlights the potential of MARL methods for scalable and robust LLM collaboration and encourages future exploration of more cooperation schemes in large-scale software systems.
7 Acknowledgments
@@
This work was partially funded by U.S. National Science Foundation (NSF) grants #2044993 and #2409351. This work used Delta and DeltaAI advanced computing and data resources at the National Center for Supercomputing Applications through allocation CIS250443 and CIS250554 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support program, which is supported by NSF grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@
We thank Gregory Bauer and Brett Bode for their help with resolving job failure issues on Delta GPUs, and members of the Lab for Learning and Planning in Robotics (LLRP) for valuable discussion.
Appendix A Formalization of Multi-Agent Interaction
@@
Many studies adopt Partially Observable Stochastic Games (POSG) to model the LLM interaction in MAS Slumbers et al. [2024], Park et al. [2025], Liu et al. [2025a], Sarkar et al. [2025]. In this section, we show that Dec-POMDP offers special merits compared to POSG in the solution concept in the cooperative settings, thus more suited to model LLM collaboration.
A.1 Dec-POMDP
@@
A Dec-POMDP is defined by $\langle\mathcal{I},\mathcal{S},\{\mathcal{O}_{i}\},\{\mathcal{A}_{i}\},R,T,H\rangle$. At each step $t$, since an agent cannot directly observe the state $s_{t}$, it usually maintains local observation-action history $h_{i,t}=(o_{i,0},a_{i,0},\ldots,o_{i,t})$ to infer a belief over the underlying state. Decisions are made according to a local policy $\pi_{i}:\mathcal{H}_{i,t}\rightarrow\Delta(\mathcal{A}_{i})$, which maps histories to probability distributions over actions. The set of all local policies forms the joint policy $\boldsymbol{\pi}=\{\pi_{1},\ldots,\pi_{n}\}$. In cooperative settings, the objective is to maximize shared cumulative rewards. As proved in Oliehoek et al. [2008], there is always an optimal joint policy in a Dec-POMDP,
@@ A.2 POSG
A.3 Non-Optimality of POSG Solutions
@@
Appendix B Cooperation Schemes

### v3

@@ Introduction
We adopt a single model and 3 multi-agent methods as our baselines. To minimize the influence of prompts on our comparison, we keep the task description fixed and only add minimal instructions. In the single-agent baseline, we prompt with the article, the agent’s role (summarizer or expanding writer), and specific format instructions. Naive concatenation builds on it by dividing the task into subtasks, assigning each agent a specific portion to complete in parallel without explicit communication. The sequential pipeline introduces one-way communication, allowing one agent to respond based on both the task description and the other agent’s output. The one-round discussion baseline enables bidirectional communication: agents first receive the same prompts as in naive concatenation, then the prompts are augmented with the other’s first-turn response in the second turn. All baseline methods operate without fine-tuning and depend solely on prompt-level interactions. Detailed prompts for each baseline are in the Appendix.
(a) Single-Turn MAGRPO on HE
(b) Multi-Turn MAGRPO on HE
(c) Single-Turn MAGRPO on CHE
(d) Multi-Turn MAGRPO on CHE
Figure 3: Normalized returns on coding collaboration: (a) structural wellness (dashed grey); (b) syntax correctness (dashed green); (c) Test score (dashed red); (d) cooperation rewards (dashed yellow); (e) total return (solid blue).
Method Dataset Efficiency Code Quality (%) Return (%)
Speed Response Time Structure Syntax Tests Cooperation
Single Model HE 73.1 1.6 100.0 100.0 64.8
CHE 65.5 1.4 100.0 100.0 63.4
Fine-Tuned Single Model HE 72.2 1.6 100.0 99.2 65.6
CHE 60.6 1.4 100.0 100.0 66.7
Naive Concatenation HE 194.9 1.1 96.1 90.6 42.5 22.7 53.9
CHE 189.4 1.1 97.5 95.0 40.1 24.0 54.3
Sequential Pipeline HE 99.6 2.2 98.4 96.5 56.4 35.1 63.1
CHE 97.4 2.0 97.5 96.3 55.2 35.2 62.5
One-Round Discussion HE 82.5 2.8 98.1 94.8 41.2 30.2 57.5
CHE 78.3 2.8 97.5 96.3 41.9 34.8 59.5
Single-Turn MAGRPO (Ours) HE 190.0 1.5 100.0 97.8 61.6 83.4 83.7
CHE 192.4 1.5 98.8 97.5 71.2 83.7 86.0
Multi-Turn MAGRPO (Ours) HE 96.5 2.7 99.7 96.9 68.4 84.9 86.7
CHE 97.1 2.5 98.6 98.2 75.0 86.3 88.5
@@
Table 2: Performance comparison of MAGRPO against baselines on HE and CHE. Speed (tokens/s) and response time (s) are recorded on GeForce RTX 5090s. Results are normalized to the return scale and averaged over 10 runs; rewards are level-based. Bolds indicate the best performance of each metric on each dataset.
Results
In this experiment, we apply MAGRPO to optimize the dual Qwen3-1.7B system in one turn. Figure 2(a) and Figure 2(b) show the evaluation results on TLDR and arXiv over 10 runs. The upward trend on all metric curves indicates that 2 agents gradually cooperate to generate coherent and consistent content with a well-organized structure. In the TLDR summarization, while the structure and logical coherence monotonically increase throughout training, the style consistency curves exhibit a decrease in the first 100 steps. This occurs as agents temporally diverge in styles to optimize other cooperative objectives, but their styles are gradually realigned and stabilized with sufficient training.
As shown in Table 1, MAGRPO is 3 times faster compared to the single Qwen3-4B model, which has a comparable number of parameters to our dual Qwen3-1.7B system. Despite receiving detailed instructions, Qwen3-4B fails to produce well-structured responses. A similar issue appears in TLDR summarization but not in arXiv expansion under multi-agent settings. This is because the outputs of homogeneous agents are naturally similar in length, which fortuitously aligns with the preference of the reward model.
Among the multi-agent baselines, parallel generation is the only one that achieves a comparable speed to ours, but it fails to generate well-structured and coherent texts due to the lack of cooperation. Sequential generation and discussion-based approaches occasionally enhance coordination through specific prompts. However, they still underperform ours in efficiency and coherence, resulting in lower total return. The limited effectiveness of prompt-instructed coordination constrains their scalability to more complex scenarios involving large numbers of agents or extended multi-turn interactions (debatefail).
Coding Collaboration
In large-scale software development, numerous developers collaborate to implement complex systems. Employing LLMs as developers is a promising direction, but coordinating them is challenging due to diverse cooperation schemes and complex failure modes. We simplify this task by using 2 Qwen2.5-Coder-3B agents to generate Python functions collaboratively. A helper agent produces auxiliary functions to support a main function generator, without any direct communication. The outputs from both agents, along with required libraries, are aggregated into complete code snippets.
HumanEval
We evaluate MAGRPO on the HumanEval (HE) dataset, which contains 164 handwritten programming problems, each containing a natural language description (prompt), a function signature (entry_point), and a set of unit tests (test). To guide learning, we design a level-based reward model that prioritizes fundamental aspects of code generation. Structural integrity verifies the presence and correctness of both main and auxiliary function definitions; syntactic correctness ensures compliance with Python syntax; test pass rate assesses functional correctness based on the proportion of successfully passed unit tests; and a cooperation quality bonus is granted when the main function properly invokes and utilizes the auxiliary function. Rewards are accumulated only when all requirements at each preceding level are satisfied.
CoopHumanEval
Some entries in HumanEval (HE) are not designed for coding collaboration; certain atomic operations (e.g., strlen(string)) can hardly be decomposed in a way that has meaningful cooperation. These noisy instances bias it towards invalid cooperation, such as merely wrapping the auxiliary function, thereby making training unstable. Thus, we construct a cooperative code generation dataset, CoopHumanEval (CHE), which comprises both original HE problems with cooperative potential (e.g., prime_fib(n)) and additional problems (e.g., unique_digits(x)). CHE problems are decomposable, where agents can explore effective cooperation schemes.
Baselines
We adopt the fixed and fine-tuned single model, and 3 multi-agent methods on fixed base models, as our baselines. In the single-agent setting, the Qwen2.5-Coder-7B model generates a Python function based on the problem description in prompt. We also fine-tune this model on the training set to adapt it to this task. In the multi-agent setting, 2 Qwen2.5-Coder-3B models serve as agents: one generates a helper function, and the other produces the main. To minimize the influence of prompts on our comparison, we keep the problem description fixed and only add minimal coordination instructions. In the naive concatenation, agents are informed of their roles and generate outputs in parallel without communication. The sequential pipeline allows the main agent to respond based on the other’s output. In the one-round discussion, agents first receive the same prompts as naive concatenation, then the prompts are augmented with the other’s last-turn response in the subsequent turns.
Results
We optimize the interaction between 2 agents using single-turn and multi-turn MAGRPO. In the multi-turn setting, problem descriptions and the agents’ initial responses are provided to static checking (AST) and dynamic execution models, which report errors for each agent. Ablations of external tools used, including self-evolving and expert guidance, are provided in the Appendix.
Figure 3(a) and Figure 3(b) show the normalized return of MAGRPO on HE over 10 runs. Single-turn MAGRPO training improves the syntactical correctness and develops valid cooperation, while the test pass rate does not show much progress. As for the multi-turning training, agents are initially overwhelmed by the external model’s feedback, resulting in even lower initial returns. They gradually adopt the error signals and improve their returns. However, the test pass rate shows no significant improvement over the single-agent model, due to noisy entries in the dataset and hence unreliable feedback. This reflects the complexity and delicacy of decentralized coder coordination, where the main agent must accurately infer the functionality of auxiliary modules and trust their correctness without communication.
The performance of single-turn and multi-turn MAGRPO on the CHE dataset is shown in Figures 3(c) and 3(d). Results show that MAGRPO achieves higher overall returns and lower variance when trained on CHE over HE. In the multi-turn setting, although agents initially struggle to interpret the feedback, like training on HE, the returns gradually improve and eventually surpass those in the single-turn training. This demonstrates that, when trained on a dataset with well-defined cooperative structures, agents can learn to utilize error messages to improve their response quality.
@@ Results
Table 2 presents a performance comparison between MAGRPO and baselines on both HE and CHE. By GRPO fine-tuning, the performance of Qwen2.5-Coder-7B model only improves slightly as the logic of test problems differs substantially. Compared to a single model, the naive concatenation method has lower test pass rates, as the main agent may generate code based on incorrect assumptions about auxiliary functionality. In the sequential pipeline method, the main agent can provide a backup for the auxiliary function when it identifies potential vulnerabilities, thereby improving the test pass rate. However, this comes at the cost of slower inference speed. Although the one-round discussion method involves more communication between agents, its effectiveness remains limited because the agents’ mutual adaptations to each other’s last responses can become misaligned. MAGRPO outperforms all baselines on both CHE and HE by facilitating effective cooperation and leveraging feedback from the external model. Additional results, including pass@k, are presented in the Appendix.
Cooperation Schemes
@@
MAGRPO identifies diverse cooperation schemes. In some cases, the auxiliary function handles the core logic, while the main agent adds backup logic or decorations to improve the overall solution. Alternatively, the main agent may act as a coordinator, decomposing the problem and assigning subtasks to the auxiliary agent. The auxiliary function may serve as a strategy filter, guiding the main agent to generate code for specific cases. While coordinator and strategy-filter schemes can improve inference efficiency, they are more prone to syntax and logical errors. With limited cooperation-oriented training data, the main agent typically resorts to more conservative roles, i.e., fallback or decoration. These cooperation schemes emerge during training under a relatively simple joint reward. More refined design patterns will likely be found when training agents to develop large-scale coding projects.
Conclusion
@@
In this paper, we model LLM collaboration as a cooperative MARL problem and formalize it as a Dec-POMDP. We propose the MAGRPO algorithm to optimize agent cooperation through shared rewards. Our experiments in coding and writing collaboration show that MAGRPO enables agents to efficiently generate high-quality responses via effective collaboration. Our work encourages future exploration of MARL-based methods for scalable and robust LLM collaboration.
Acknowledgments
@@
This work was partially funded by NSF grants #2044993 and #2409351. It used Delta and DeltaAI computing resources at the National Center for Supercomputing Applications through allocation CIS250443 and CIS250554 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support program, which is supported by NSF grants #2138259, #2138286, #2138307, #2137603, and #2138296.
@@
We thank Tianle Chen for improving our MARL training framework for LLM collaboration, CoMLRL, and members of the Lab for Learning and Planning in Robotics (LLPR) for the valuable discussion. We thank Gregory Bauer and Brett Bode for helping us resolve job failure issues.
Appendix A Formalizations of Multi-Agent Interaction
@@
Many studies adopt Partially Observable Stochastic Games (POSG) to model the LLM interaction in MAS (fama; spiral; amongus; marti2025). In this section, we show that Dec-POMDP offers special merits compared to POSG in the solution concept in the cooperative settings, thus more suited to model LLM collaboration.
Dec-POMDP
@@
A Dec-POMDP is defined by $\langle\mathcal{I},\mathcal{S},\{\mathcal{O}_{i}\},\{\mathcal{A}_{i}\},R,T,H\rangle$. At each step $t$, since an agent cannot directly observe the state $s_{t}$, it usually maintains local observation-action history $h_{i,t}=(o_{i,0},a_{i,0},\ldots,o_{i,t})$ to infer a belief over the underlying state. Decisions are made according to a local policy $\pi_{i}:\mathcal{H}_{i,t}\rightarrow\Delta(\mathcal{A}_{i})$, which maps histories to probability distributions over actions. The set of all local policies forms the joint policy $\boldsymbol{\pi}=\{\pi_{1},\ldots,\pi_{n}\}$. In cooperative settings, the objective is to maximize shared cumulative rewards. As proved in (fransoptimalq), there is always an optimal joint policy in a Dec-POMDP,
@@
Non-Optimality of POSG Solutions
@@
Method Dataset Pass@k (%) Acc@k (%) Coop@k (%)
@3 @5 @10 @3 @5 @10 @3 @5 @10
Single Model HE 67.7 71.0 83.9 85.4 87.9 95.1
CHE 68.8 75.0 81.3 75.0 81.3 88.8
Fine-Tuned Single Model HE 69.2 72.9 84.6 83.0 89.2 95.6
CHE 69.4 74.5 83.7 74.3 82.2 90.3
Naive Concatenation HE 45.2 51.6 64.5 70.4 75.5 80.9 49.5 67.7 76.3
CHE 43.8 56.3 68.8 57.0 63.8 73.8 47.9 69.3 81.3
Sequential Pipeline HE 56.8 61.3 71.0 78.8 84.5 91.9 62.4 73.3 92.7
CHE 75.0 81.5 87.5 88.2 89.5 91.3 75.0 75.0 81.3
One-Round Discussion HE 51.6 61.3 71.0 71.8 81.3 87.5 58.1 70.0 78.7
CHE 50.0 68.8 68.8 75.4 82.5 82.0 66.7 68.7 75.0
Single-Turn MAGRPO (Ours) HE 64.8 68.1 71.0 75.3 76.3 86.4 83.8 90.3 90.3
CHE 75.0 75.0 81.2 80.0 82.5 87.5 87.5 93.8 93.8
Multi-Turn MAGRPO (Ours) HE 71.0 80.6 90.3 85.7 92.6 94.7 93.5 96.8 96.8
CHE 75.0 81.5 87.5 86.4 92.5 95.4 93.8 96.8 100.0
Table 5: Performance comparison between MAGRPO and baseline methods with pass@k, acc@k, coop@k, on HE and CHE. The bold texts indicate the best performance of each metric on each dataset.
@@
Appendix B Additional Results
We report additional results in this section to validate the effectiveness of our approach.
@k Ablation
In LLMs, single-run inference often leads to high variance. To provide a more reliable evaluation of the capacities (yue2025doesreinforcementlearningreally), we evaluate the pass, test accuracy, and cooperation at $k$ runs (pass@k, acc@k, and coop@k) on MAGRPO and baselines. These @k metrics measure the best outcome among $k$ generated solutions for each problem and are averaged over all problems in $\mathcal{D}$.
Pass@k is calculated as the probability that at least one out of $k$ generated solutions passes all test cases (evaluate). Specifically, a set $\mathcal{P}$ of $k$ solutions is randomly sampled from a pool of $\mathcal{M}$ generated solutions. To make it consistent with other @k metrics, we express pass@k as,
$\text{Pass@k}=\frac{1}{|\mathcal{D}|}\sum_{j=1}^{|\mathcal{D}|}\mathbb{E}_{\mathcal{P}\sim\text{Sample}(\mathcal{M},k)}\left[\max_{p\in\mathcal{P}}\mathds{1}(n_{j}^{\text{pass}}=n_{j})\right],$ (6)
where $\mathds{1}(n_{j}^{\text{pass}}=n_{j})$ is the indicator function that equals 1 if all test cases are passed in problem $j$ (i.e., the number of passed tests $n_{j}^{\text{pass}}$ equals the total number of tests $n_{j}$), and 0 otherwise.
However, pass@k does not offer a fine-grained assessment of solution quality, as functional correctness is represented as a binary variable. As a result, failing a single test case is treated the same as failing all in pass@k. To provide a more unbiased evaluation, we use accuracy@k, defined as the highest test accuracy among the $k$ generated solutions, averaged across all problems.
$\text{Acc@k}=\frac{1}{|\mathcal{D}|}\sum_{j=1}^{|\mathcal{D}|}\mathbb{E}_{\mathcal{P}\sim\text{Sample}(\mathcal{M},k)}\left[\max_{p\in\mathcal{P}}\frac{n_{j}^{\text{pass}}}{n_{j}}\right],$ (7)
where $n_{j}^{\text{pass}}$ and $n_{j}$ are the number of passed and total unit tests of problem $j$, and $n_{j}^{\text{pass}}/n_{j}$ is the test accuracy.
Similar to acc@k, we also propose coop@k, which measures the average of the highest normalized cooperation return achieved among $k$ runs across all problems. Formally, the Coop@k is defined as,
$\text{Coop@k}=\frac{1}{|\mathcal{D}|}\sum_{j=1}^{|\mathcal{D}|}\mathbb{E}_{\mathcal{P}\sim\text{Sample}(\mathcal{M},k)}\left[\max_{p\in\mathcal{P}}R^{\text{coop}}_{p}\right],$ (8)
where $R_{p}^{\text{coop}}=\sum_{t=0}^{H-1}r^{\text{coop}}_{p,t}$ is the cooperation return over horizon $H$, and $r^{\text{coop}}_{p,t}$ denotes the cooperation reward obtained by solution $p$ at turn $t$.
We generate $15$ samples in $\mathcal{M}$ and evaluate all methods with $k=10$. Table 5 presents the results for pass@k, acc@k, and coop@k at $k=3,5,10$, comparing MAGRPO with baseline methods. As expected, the trends in @k metrics are consistent with the @1 results. By fine-tuning with GRPO, the pass@k, acc@k, and coop@k are slightly improved. The naive concatenation method remains worse than the single-agent baseline in terms of pass@k, as the main agent may generate code based on incorrect assumptions about the auxiliary function. The sequential pipeline mitigates this issue by allowing the main agent to reference the auxiliary output during generation, yielding substantial improvements across all @k metrics, particularly on CHE. Although the one-round discussion method shares the same prompts as naive concatenation in the first turn, the additional discussion round yields limited improvement across the @k metrics due to misalignment in cross-adaptation.
Multi-turn MAGRPO outperforms all baselines across most @k metrics by leveraging feedback from external sources. These more comprehensive experiments further validate the general effectiveness of our approach and highlight promising directions for future work, such as training LLMs through interactions with static analyzers, sandbox-based evaluations, or expert models. Notably, the acc@k metric offers a more fine-grained perspective on performance trends as $k$ varies. In some cases, slightly increasing $k$ may not yield additional solutions that pass all test cases (as measured by pass@k), yet the improvement is still captured through higher accuracy.
Cooperation Schemes